OpenStudy (anonymous):

A system of equations is shown below: 6x – 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

I know the answer.

OpenStudy (anonymous):

... Well?

OpenStudy (anonymous):

Eggplant.

OpenStudy (misssunshinexxoxo):

Ok please stop spamming the question kamui

OpenStudy (misssunshinexxoxo):

Hello; are you familiarized on how to solve system of equations?

OpenStudy (anonymous):

Gee thanks man! I'm totally going to pass this test. @kamuigakupo And no not really @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (anonymous):

Eh?

OpenStudy (misssunshinexxoxo):

That is your answer

OpenStudy (misssunshinexxoxo):

11x+y=7 6x-2y=3 5x+3y=4

OpenStudy (anonymous):

Oh I see! Thanks! Do you think you can help me with another?

OpenStudy (misssunshinexxoxo):

Please close this question and make a new one tag me:)